Find an equation of the ellipse having a major axis of length 10 and foci at , (1,−2 )and , (−7,−2.)


1
Expert's answer
2022-02-22T13:42:51-0500

The equation of an ellipse is

"\frac{\left(x - h\right)^{2}}{a^{2}} + \frac{\left(y - k\right)^{2}}{b^{2}} = 1,"

where "\left(h, k\right)" is the center, "a" and "b" are the lengths of the semi-major and the semi-minor axes.

Thus "a=10\/2=5."

From properties of an ellipse


"k=-2"

"(h-1)^2=a^2-b^2"

"(h+7)^2=a^2-b^2"

Then


"(h-1)^2=(h+7)^2"

"h^2-2h+1=h^2+14h+49"

"16h=-48"

"h=-3"

"(-3-1)^2=5^2-b^2"

"b^2=9"

The equation of the ellipse is


"\frac{\left(x+3\right)^{2}}{25} + \frac{\left(y +2\right)^{2}}{9} = 1,"